OpenStudy (anonymous):

Which of the following equations is the slope-intercept form of 4x - 2y = 8? A. y = 2x - 4 B. y = -4x - 8 C. y = 1/2 x + 2 D. y = -2x + 4

OpenStudy (anonymous):

b- y=-4x-8

OpenStudy (anonymous):

I think...

OpenStudy (hoblos):

general equation in slope-intercept form is y=mx+b where m is slope b is y-intercept 4x - 2y = 8 2y = 4x-8 y=2x+4

OpenStudy (anonymous):

its a

OpenStudy (hoblos):

yeah y = 2x - 4 not y = 2x + 4 a typo though im sorry

OpenStudy (anonymous):

oooohh.. yeah I agree with HOBLOS.. lol can you help me with my first question now?..